OpenStudy (anonymous):

Approximate the solution to 8x^3+12x^2+6x+1=4

OpenStudy (anonymous):

8x^3 + 12x^2 + 6x = 3 4x^3 + 6x^2 + 3x = 3

OpenStudy (anonymous):

8x^3 + 12x^2 + 6x -3 = 0 Here, one method is to use Trial and Hit Method and then by using Long Division, we can find the factors for this Cubic polynomial and hence we can find the value of x..